ms. Jones invested 18,000 in two accounts. One account pays 6 % simple interest and the other pays 8%. Her total interest for the year was $1,290. How much did she have in each account



can anyone help me find the answer with a full break down

Answer by stanbon(48558)

ms. Jones invested 18,000 in two accounts. One account pays 6 % simple interest and the other pays 8%. Her total interest for the year was $1,290. How much did she have in each account
-------------------------
Equation:
Quantity: x + y = 18000
Interest:0.06x + 0.08y = 1290
--------------------------------------
Multiply the Quantity Eq. by 6
Multiply the Interest Eq. by 100
6x + 6y = 108000
6x + 8y = 129000
----------------------------------------
Subtract the 1st equation from the 2nd and solve for "y":
2y = 21000
y = 10500 (amount invested at 8%)
---------------------
Since x+y= 180000, x = 7500 (amount invested at 6%)
====================================
Cheers,
Stan H.